It is known that 2.6% of the population has a certain disease. A new test is developed to screen for the disease. A study has shown that the test returns a positive result for 18% of all individuals, and returns a positive result for 92% of individuals who do have the disease. If a person tests positively for the disease under this test, what is the probability that they actually have the disease? 0.1276 0.1329 0.1435 0.1223 O 0.1382
Answer #1

Answer

Back-up Theory

If A and B are two events such that probability of B is influenced by occurrence or otherwise of A, then Conditional Probability of B given A, denoted by P(B/A) = P(B ∩ A)/P(A)…….....................................................….(1)

P(A/B) = P(B/A) x {P(A)/P(B)}……………………………..……………............................……..………….(2)

Now to work out the solution,

Let A represent the event that an individual has the disease and B represent the event

that the test returns positive.

Trivially, AC represents the event that an individual does not have the disease and

BC represents the event that the test returns negative.

2.6% of the population has the disease = > P(A) = 0.026 …………………………………..........…..(3)

Test returns positive in 18% of cases = > P(B) = 0.18….. ……………………………….........……..(4)

Test returns positive in 92% of individuals who do have the disease

= > P(B/A) = 0.92….. ……………………………………………………………...............……….…..(5)

We want to find the probability that a person has the disease, if the person tests positive. i.e.,

P(A/B)

= P(B/A) x P(A)/P(B) [vide (2)]

= (0.92 x 0.026)/0.18 [vide (5), (3) and (4)]

= 0.1329 Second Option Answer

DONE